OpenStudy (anonymous):

Prove the explicit formula for the sequence defined recursively by \[a _{1}=1\] and \[a _{k}=2a_{k-1}+6\] is \[a _{n}=7(2^{n-1})-6\] (use mathematical induction)

OpenStudy (anonymous):

Here's where i'm at thus far:

OpenStudy (anonymous):

let p(k) be: \[a _{k}=7(2^{n-1})-6\] 1. show p(1) is true (that a_1=1): \[a _{1}=7(2^{(1)-1})-6 = 7(1)-6=1\]

OpenStudy (mr.math):

Isn't this the same problem we did last time?

OpenStudy (anonymous):

2. show p(k) implies p(k+1) p(k+1): \[a _{(k+1)}=7(2^{(k+1)-1})-6=7(2^{k})-6\]

OpenStudy (anonymous):

We found the formula yes

OpenStudy (anonymous):

I am attempting to prove it with induction now

OpenStudy (mr.math):

I see!

OpenStudy (anonymous):

@mr math i always forget how to do these, and have to look it up each time (not the proof by induction, finding the formula) can you link to this answer, i would love to see it

OpenStudy (anonymous):

Yea, i can, i asked it :)

OpenStudy (anonymous):

thank. i always reinvent the wheel every time. can you complete the inductive step?

OpenStudy (anonymous):

Would I just multiply bu (1/2)?

OpenStudy (anonymous):

that would give me 7(2^k-1) like i want in p(k)

OpenStudy (anonymous):

but would also make 6, 3 which we dont want...

OpenStudy (anonymous):

but thats not right, I need to modify p(k) to look like p(k+1), not the other way around

OpenStudy (anonymous):

so multiply p(k) by 2.. would be the same issue, 6 as 12

OpenStudy (mr.math):

You need to prove that for every integer \(k\ge1\) the explicit formula gives the same value of \(a_k\) as the recursive formula. This is your aim.

OpenStudy (mr.math):

You showed that for n=1. Now prove that if it's true for n=k then it's also true for n=k+1.

OpenStudy (anonymous):

Right, and I show that by modifying one in such a way that they yield the same result

OpenStudy (anonymous):

\[a _{k}=2a_{k-1}+6\] by definition, so your job is to prove that \[a_{k+1}=7(2^{k})-6\] GIVEN that \[a_k=7(2^{k-1})-6\]

OpenStudy (anonymous):

use first the definition, and then the inductive hypothesis, and it should work out swimmingly

OpenStudy (mr.math):

satellite explained it in the best way I could possibly think of.

OpenStudy (anonymous):

Yes, so I need to do some algebra on the 2nd equation

OpenStudy (mr.math):

Just one step actually.

OpenStudy (anonymous):

All I can see to do is multiply by 2, but that changes the 6..

OpenStudy (mr.math):

Start from the definition with n=k+1 and use the assumption about n=k to get the explicit formula for n=k+1.

OpenStudy (mr.math):

What's \(a_{k+1}\) by definition?

OpenStudy (anonymous):

not 7(2^k)-6

OpenStudy (mr.math):

\(a_{k+1}=2a_k+6\), right?

OpenStudy (anonymous):

Ok yea

OpenStudy (mr.math):

But what's \(a_k\) based on your assumption?

OpenStudy (anonymous):

I was trying to add something to that expression with a_k-1

OpenStudy (mr.math):

Not following.

OpenStudy (anonymous):

\[a _{k}=7(k ^{k-1})-6\] base on our assumption

OpenStudy (mr.math):

Plug this into the last expression I gave.

OpenStudy (anonymous):

comes to \[7(2^{k})\]

OpenStudy (mr.math):

Hmm, We have "by definition" \[a_{k+1}=2a_k+6\] But we've assumed that \(a_k=7(2^{k-1})-6\), thus \[a_{k+1}=2(7(2^{k-1})-6)+6=7(2^k)-6.\]

OpenStudy (mr.math):

Is this the same as what we had in the explicit formula? YES! PROVED!

OpenStudy (mr.math):

You can see it, right?

OpenStudy (mr.math):

Apparently you didn't multiply -6 by 2, that's why you got a wrong expression.

OpenStudy (anonymous):

Ah, ok

OpenStudy (anonymous):

So we dont even need p(k+1) version of the explicit formula?

OpenStudy (anonymous):

just a_k+1 and p(k)?

OpenStudy (anonymous):

This isnt the same as our explicit formula.

OpenStudy (anonymous):

All it is is our p(k+1) formula

OpenStudy (mr.math):

Of course it's the same as our explicit formula evaluated at \(n=k+1.\)

OpenStudy (mr.math):

Check by plugging n=k+1 in the formula and check for yourself :D

OpenStudy (anonymous):

http://imgur.com/Of0yw

OpenStudy (anonymous):

What i've written doesn't seem right to me for some reason

OpenStudy (mr.math):

It's confusing me too. I think what we wrote above is pretty clear. You just need to read it again. Always remember we want to show that OUR explicit formula gives the same values that the original definition would give.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

Whats confusing is usually with these we show p(k) can be shown to be equivalent to p(k+1) using algebra of some sort, and that is sufficient proof

OpenStudy (mr.math):

We started from the definition, used our assumption about a_k to and showed that it's equal to the explicit formula we derived.

OpenStudy (anonymous):

Here, it seems we're showing an alternate definition of p(k+1) can be made to look as the other definition of p p(k+1) using p(k)

OpenStudy (mr.math):

We're using the definition you wrote in your question.

OpenStudy (anonymous):

yea, the definition is just an alternate version of p(k+1)

OpenStudy (mr.math):

What's p(k+1)?

OpenStudy (mr.math):

7(2^k)-6?

OpenStudy (anonymous):

p(k+1): \[a _{(k+1)}=7(2^{(k+1)-1})-6\]

OpenStudy (anonymous):

yes

OpenStudy (mr.math):

That's actually what wanted to show (i.e the statement p(k+1) is true).

OpenStudy (anonymous):

and a_k+1 also has the alternate meaning if we use the original definition, (not p(k) or p(k+1)

OpenStudy (mr.math):

@satellite: Help me out here, you probably have better explaining abilities :D

OpenStudy (anonymous):

ok i will give it a shot but no guarantee (thank you for the compliment though)

OpenStudy (anonymous):

our overall goal is to show the explicit formula is true

OpenStudy (anonymous):

in the 2nd step of induction our goal is to show that p(k) "implies" p(k+1)

OpenStudy (anonymous):

it is my understanding that we can show this implication by showing p(k) can be modified algebraically to look the same as p(k+1)

OpenStudy (anonymous):

we have the following definition \[a _{k}=2a_{k-1}+6\]right? and through some other method (certainly not induction) we found an explicit formula i.e. a formula for \[a_k\] that did not depend on the recursion now we wish to show that it works for all k

OpenStudy (anonymous):

so we are going to assume that this formula \[a _{k}=7(k ^{k-1})-6\]works for k, and show that it also works for k + 1

OpenStudy (anonymous):

in other words we know that \[a _{k}=7(k ^{k-1})-6\] and now we want to show that \[a _{k+1}=7(k ^{k})-6\] as well

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

are you meaning to write 2^k-1

OpenStudy (anonymous):

not k^k-1

OpenStudy (anonymous):

yeah that is a typo

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i follow thus far

OpenStudy (anonymous):

we get to assume \[a _{k}=7(2 ^{k-1})-6\] and we want to show that \[a _{k+1}=7(2^{k})-6\]

OpenStudy (anonymous):

but we have a definition for \[a_{k+1}\]

OpenStudy (anonymous):

it is \[a_{k+1}=2a_k+6\]

OpenStudy (anonymous):

and now we assume (by induction) that our formula works for k, that is we know that \[a _{k}=7(2 ^{k-1})-6\] so where we see an \[a_k\]in the above recursion we get to replace it by \[a _{k}=7(2 ^{k-1})-6\]

OpenStudy (anonymous):

ok, so, in induction, it can both ways to prove the implication?

OpenStudy (anonymous):

here we go... \[a_{k+1}=2a_k+6=2\times [7(2^{k-1})-6]+6\]

OpenStudy (anonymous):

now we do some algebra on the second part, and if we get \[a_{k+1}=7(2^{k})-6\] then we win, because that is the formula we are trying to prove

OpenStudy (anonymous):

Ok so i guess this is just a different method of induction I am used to

OpenStudy (anonymous):

In my previous inductions we had to prove the implication, that is, show p(k) could be modified in some way to resemble p(k+1)

OpenStudy (anonymous):

here we are just using both p(k) and p(k+1) to get the explicit formula

OpenStudy (anonymous):

i am fairly sure that it may look different but is the same. btw you notice that induction in a method of proof, but does not help a whit in finding the formula that is entirely different. yes, that is exactly what we did. look carefully. i think i see the confusion, because we had to work form the definition to get to the inductive step. in real life the hard part of a proof by induction is reducing to the previous step

OpenStudy (anonymous):

so showing p(k+1) can be modified (using p(k) which we know is true) to resemble our explicit formula is proof that p(k+1) is true

OpenStudy (anonymous):

is it still appropriate say "show p(k) => p(k+1)" for the 2nd step?

OpenStudy (anonymous):

hmm, that pdf isnt working for me

OpenStudy (anonymous):

You two here?